MG929997:

Jamilia has to purchase a new car and is putting 1/3 of the price down. She then finances the rest at a 1% interest rate . If she ended up paying $12000 after paying the 1/3 down how much did the car cost (round to the nearest dollar )

MrMudd183:

Can you post the answer choices please?

supie:

Ok so we would have to find what \[1\%\ of\ 12,000\ =\] since 12,000 is the total price, So what we would have to do to find that is \[1/100(12000)=\]\[(1/100)(12000/1)=\]\[(1)(12000)/(100)(1)=\]\[12000/100=120\] So then we would have to do \[12000-120=?\] I think....